# Yanjie Pan1∗    Qingdong He2∗    Zhengkai Jiang3    Pengcheng Xu4    Chaoyi Wang5    Jinlong Peng2 Haoxuan Wang1    Yun Cao2    Zhenye Gan2    Mingmin Chi1    Bo Peng1    Yabiao Wang2

### v1

@@ Abstract
Recent advances in diffusion-based text-to-image generation have demonstrated promising results through visual condition control. However, existing ControlNet-like methods struggle with compositional visual conditioning - simultaneously preserving semantic fidelity across multiple heterogeneous control signals while maintaining high visual quality, where they employ separate control branches that often introduce conflicting guidance during the denoising process, leading to structural distortions and artifacts in generated images. To address this issue, we present PixelPonder, a novel unified control framework, which allows for effective control of multiple visual conditions under a single control structure. Specifically, we design a patch-level adaptive condition selection mechanism that dynamically prioritizes spatially relevant control signals at the sub-region level, enabling precise local guidance without global interference. Additionally, a time-aware control injection scheme is deployed to modulate condition influence according to denoising timesteps, progressively transitioning from structural preservation to texture refinement and fully utilizing the control information from different categories to promote more harmonious image generation. Extensive experiments demonstrate that PixelPonder surpasses previous methods across different benchmark datasets, showing superior improvement in spatial alignment accuracy while maintaining high textual semantic consistency.
@@
Figure 1: Visualization comparison between Uni-ControlNet [51] and our proposed method in different conditional controls with the same text prompt. (a, left two columns) Text and various visual controls, where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ denotes edge, sketch, depth and pose map respectively. (b, middle three or four columns) Generation results from UniControlNet. (c, last column) Generation results from our PixelPonder. Previous methods struggled to generate coherent results under multiple conditions, while our results maintain strong similarity to the respective visual controls.
footnotetext: Equal contribution     Corresponding author
1 Introduction
@@
The emergence of diffusion-based text-to-image synthesis [32, 9, 25, 26, 29, 30] have driven significant progress in single-modality conditioned generation frameworks [48, 20, 22, 4, 21]. While these approaches demonstrate enhanced controllability through visual-textual alignment, fundamental limitations persist due to modality-specific biases in visual conditioning. As evidenced in [35, 11], distinct visual control signals exhibit complementary yet conflicting characteristics, such as depth maps effectively regulate inter-object spatial relationships but lack fine-grained object details, whereas canny maps capture precise texture contours while disregarding global structural context. This modality specialization gap raises a critical challenge: How to establish an effective fusion paradigm for multi-visual condition control that synergistically integrates heterogeneous visual features (e.g., structural priors, textural details) while resolving inter-modality conflicts?
In prior research endeavors, numerous advanced methods for single visual condition control [49, 4, 23] have been developed, yielding promising results in controllable image generation. Building upon this foundation, subsequent studies [51, 27, 22] have investigated unified approaches for single visual condition control generation. These methods successfully integrate various image condition controls into a single model, thereby accomplishing the goal of multi-category visual independent control generation using a unified framework. Despite these advancements, there have been few efforts to extend these methods to jointly control image generation with multiple visual conditions, yielding some positive outcomes. Furthermore, the prevalent strategy of leveraging prior knowledge from the control branches of each visual condition for simple combinations of visual conditions presents significant challenges. As illustrated in Fig. 1, generating normal images under complex, highly coupled visual condition controls remains a difficult task.
@@
In practical applications, it is often necessary to describe the visual features of a significant object using multiple visual conditions to achieve precise control over its generation. Users typically desire to simultaneously control both the layout and the intricate details of the key target object. However, visual conditions that are rich in layout and detail information are challenging to integrate into a single visual condition map. A common scenario involves using canny maps to represent details and depth maps to control the layout. In such case, these visual conditions are highly coupled, overlapping, and complementary to one another.
@@
Bridging the conflicting information among these visual conditions and reasonably utilizing the emphasis of various visual conditions to coordinate the generation of the object is exactly the issue that this paper wants to focus on. As shown in Fig. 2(a), prior efforts including Cocktail [16] and Uni-ControlNet [51] propose to consolidate various visual conditions into unified latent representations. While this coarse-grained unification alleviates inter-condition conflicts through dimensionality reduction, it remains suboptimal in preserving modality-specific attributes critical for precise control (e.g., high-frequency textures in cannys vs. low-frequency structures in depth). The inherent information bottleneck induced by latent compression disproportionately attenuates complementary cross-modal features, fundamentally limiting their capacity to synergize discriminative visual cues. This constitutes a critical limitation: current unification paradigms prioritize conflict mitigation over inter-modal feature reinforcement.
@@
Figure 2: Comparison of different schemes dealing with multiple condition. (a) Existing methods integrate various visual controls into a highly consistent visual control signal in the temporal domain through concatenation or weighted summation. (b) We employ dynamic patch selection to reconstruct visual conditions at the patch level while preserving the regional features of each visual condition in the temporal domain, enabling multi-visual controllable generation in the generative model.
@@
To address the aforementioned challenges, we introduce an innovative visual condition combination method termed PixelPonder. As illustrated in Fig. 2(b), our method distinguishes itself from previous approaches by integrating diverse visual conditions at the patch level. Recognizing that visual conditions enriched with low-frequency or high-frequency signals significantly influence various denoising stages during the image generation process [15], we have developed an adaptive patch selection mechanism that dynamically adjusts with timesteps. This mechanism decouples each visual condition in the temporal domain, enabling the use of distinct combinations of visual conditions to regulate the denoising process at different stages. Our proposed combination method amalgamates various visual conditions with greater granularity, effectively balancing the conflicts between disparate visual control information, and facilitating complementary control of high- and low-frequency information within the temporal domain. It is worth noting that when operating under a single visual condition control, our proposed method effectively reverts to the single visual control scheme. The efficacy of ControlNet’s single visual condition control has been extensively validated in numerous studies [49, 51, 27]. Therefore, this paper does not investigate the generative capabilities of our method under single visual condition scenarios. Our main contributions are summarized as follows:
@@
  • Flexible Condition Combination: Our PixelPonder supports flexible visual control combinations, regardless of the category of visual control, addressing the control conflict of multiple visual conditions for the same object.
@@
2 Related Work
2.1 Diffusion-based Generative Models
@@
Diffusion models [36, 15, 24, 26, 29, 5] have made remarkable strides in the field of image generation. Thanks to improvements in training and sampling strategies [15, 37, 38], as well as the application of latent diffusion methods [32], the computational resources required for these models to generate high-quality images have been substantially reduced, demonstrating promising prospects in various applications [4, 2, 12, 30]. In the domain of text-to-image generation, diffusion models [32] combined with U-Net [33] have successfully integrated text control into the image generation process through cross-attention mechanisms. To further enhance generative capabilities, the Transformer architecture has been integrated into diffusion models [25, 6, 10] to replace U-Net. Building on this foundation, FLUX [19], which introduces flow matching objectives, has achieved state-of-the-art performance in image generation.
@@
Figure 3: Overall pipeline of the proposed PixelPonder. The Patch Adaptaiom (Section 3.2) Module segments various visual conditions $\{V_{1},V_{2},\ldots,V_{n}\}$ into patches and dynamiclly reorganizes the visual conditions at the patch level to obtain a unified visual control signal $SP_{out}$. Subsequently, $SP_{out}$ is sent to the control network, with the Time-aware Control Injection scheme (Section 3.3) to achieve high-quality image generation under multiple visual conditions.
2.2 Controllable Text-to-Image Diffusion Models
@@
The generation of images controlled by textual descriptions often lacks fine-grained spatial control. To enhance spatial control capabilities, numerous studies [22, 49, 21, 17, 8, 46, 50] have focused on incorporating visual control signals into the generation process. In this context, ControlNet [49] proposed a network architecture consistent with the backbone network, encoding additional visual conditions into latent representations and injecting control signals into the corresponding backbone network using zero convolution, thus achieving a spatial alignment mechanism within diffusion models. Due to its straightforward approach and effective control results, it has been widely adopted in various text-to-image diffusion models [32, 19]. Furthermore, recent research has used various prompt engineering techniques [21, 46] and inter-attention constraints [7, 44] to facilitate more orderly generation.
@@
Some studies [27, 22, 42, 1] have addressed the differences and diversity of various visual conditions in different modalities using the mix of experts (MOE) to decouple different visual conditions into distinct feature subspaces, thus establishing a unified visual control architecture. Subsequently, other research has explored the possibility of generating multiple condition controls within a single diffusion model through simple feature aggregation methods [16, 51]. However, these methods are based on the low coupling of various conditions, which limits their effectiveness in tasks requiring the control of complementary conditions.
3 Method
In this paper, we introduce PixelPonder, a specialized framework for handling complex scenarios of image generation under multiple visual conditions. The pipeline of PixelPonder is demonstrated in Fig. 3. In this section, we first provide a brief overview of the previous Diffusion Transformer in Section 3.1, along with ControlNet. Then, in Section 3.2, we dive into a detailed description of the Patch Adaptation Module, which primarily focuses on the autoregressive reorganization of visual conditions at the patch level. In Section  3.3, we elaborate on the time-step awareness injection strategy and the control implementation of the control network. Finally, we summarize our training optimization objectives in Section 3.4.
3.1 Preliminary
@@
Diffusion Transformer (DiT)[41, 25] is applied in models such as FLUX  [19] and Stable Diffusion 3 [10]. Compared to the Latent Diffusion Model [32], these models utilize a transformer [34] instead of U-Net as the denoising network to iteratively process noise tokens.
@@
A model with a DiT architecture handles noisy image tokens $X\in\mathbb{R}^{N\times d}$ and text condition tokens $Y\in\mathbb{R}^{M\times d}$, where $d$ is the dimension of the embedding layer, $N$ and $M$ represent the number of tokens for images and text, respectively. In this structure, these tokens maintain consistent sizes before and after being processed by the transformer. Additionally, they redefine the mapping from images to noise using flow matching. In this rectified mapping, the data at time $t$, denoted as $z_{t}$, is transformed from the real image $x_{0}\in\mathbb{R}^{C\times H\times W}$ to pure noise $N\sim\mathcal{N}(0,I)$ at the following position:
@@
ControlNet [49] is designed to inject visual control information into a denoising network, thereby guiding the iterative process of noise tokens. Specifically, ControlNet freezes the parameters $\Theta$ of the backbone network $F(\cdot;\Theta)$ and creates a trainable copy $\Theta_{c}$ of these parameters to incorporate zero-initialized convolutional layers $Z$. When applied to generative models based on the U-Net architecture, a common practice is to add the visual condition $C_{i}$ to the noise $x_{t}$ as input to the trainable copy, while utilizing other signals $s$ that is consistent with the backbone network. This process can be formalized as follows:
@@ 3.2 Patch Adaptation
where $O_{t}$ refers to the time step information, and its specific role will be discussed in Section 3.3.
@@
1:  $SP_{0}=\textbf{0}_{N\times d}$
2:  $C_{n}=[P_{n}^{0},\ldots,P_{n}^{m}]$
3:  $\mathcal{C}_{0}=\{C_{1},C_{2},\ldots,C_{n}\}$
4:  $i\leftarrow 0$
5:  $n_{p}\leftarrow$ number of selected patches
@@
7:      // The loop of process to update $U_{i}$
8:     $M_{\text{prob}}=[\text{ISB}_{\text{C}}(\mathcal{C}_{i},O_{t},Y)]+\text{ISB}_{\text{SP}}(SP_{i})$
9:     // The patch selection process, denoted as $F_{sp}$.
10:     $M_{N,i}=\text{Zero}(\text{Top}_{r}(M_{prob}))$, $M_{sp,i}=\sim M_{N,i}$
11:     $\mathcal{C}_{i+1}=\mathcal{C}_{i}M_{N,i}$
12:     $SP_{i+1}=SP_{i}+M_{sp,i}\mathcal{C}_{i}$
13:     $i\leftarrow i+1$
14:  end while
15:  return $SP_{i}$
Assuming that the visual input to the PAM consists of $N$ types of condition images, denoted as $V_{1},V_{2},\ldots,V_{n}$, we use the corresponding visual condition feature encoders $E_{i}$ to obtain various visual condition features $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$, $C_{k}\in\mathbb{R}^{e\times d}$ and $e$ denotes the feature dimension. To analyze and process the visual condition features in a more fine-grained manner, we can consider the visual condition feature $C_{i}$ as represented by patches $P_{i}^{m}\in\mathbb{R}^{d}$, expressed as:
$C_{k}=[P_{i}^{0},\ldots,P_{i}^{m}],$ (5)
@@
Specifically, to achieve $F_{\text{sp}}$, we design the Image Stream Block (ISB). Unlike DSB, to ensure that the textual control signal remains highly consistent globally, we change the DSB text input selection strategy, abandoning the textual information $Y_{\text{DSB}}$ generated in dual-stream matching as input and instead using a globally unified textual feature representation $Y$ as the input textual information. The ISB is used to obtain the weights of each patch in various visual condition features under the current textual control $Y$ at time step $t$, denoted as $\mathcal{W}_{c}=\{W_{0,i},\ldots,W_{n,i}\}$. Furthermore, to meet requirement b) of $F_{\text{sp}}$, we apply an ISB to $SP_{i}$ outside of $\mathcal{C}_{i}$ to obtain a global weight $W_{\text{sp}}$ that represents the global composition information of $SP_{i}$ under the current $U_{i}$. The final obtained patch selection probability map $M_{\text{prob}}$ is given by:
@@
Figure 4: Denoising results at steps 1, 5, 10, 15, 20, and 25, where each pair includes the generated result and the selection results $SP_{out}$ at the current time step. In each $SP_{out}$, canny, pose, depth and sketch are indicated in red, blue, green, and yellow respectively.
Fig. 4 illustrates the visualization results of $SP_{out}$ obtained by PAM in different time steps, as well as the generated images. In the early stages of the denoising process, the generated images exhibit a noticeable tendency towards patchiness. The segmentation contours between the generated image patches are prominent and remain evident even in the later stages of denoising. Both examples in Fig. 4 demonstrate the characteristics of PAM: a) PAM exhibits a balanced tendency in its selections over various time steps. b) Notably, for patches at the same location, PAM’s global selections are not consistent; instead, they encompass patches from all categories.
@@
3.3 Time-aware Control Injection
@@ 4 Experiments
Methods Conditions FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
T2I-Adapter [22] all 66.95 71.47 24.03 57.95 64.72 74.35 31.76 55.07
Uni-ControlNet [51] all 32.58 78.08 29.37 65.85 44.35 77.40 37.98 66.75
UniControl [27] all 25.15 74.09 35.58 72.05 30.95 72.96 47.31 67.50
Cocktail [16] pose+hed 24.67 79.87 24.14 67.13 51.51 79.38 29.59 64.16
PixelPonder(Ours) all 11.85 78.60 43.99 69.54 10.61 77.12 66.60 67.32
Table 1: Comparison of generation quality, controllability, and text-image consistency on the MultiGen-20M [20] and Subject-200K [40] datasets. The best and second-best results are highlighted in red and blue, respectively. “all” under the Conditions column represents that all conditions (canny maps, sketches, pose maps, depth maps) are used for generation.
We validate the effectiveness of PixelPonder on the MultiGen-20M [20] and Subject-200K[40] datasets. Our evaluation primarily focuses on several leading methods, including T2I-Adapter [22], Uni-ControlNet [51], UniControl [27], and Cocktail [16] in the realm of multiple controllable text-to-image diffusion models. It should be noted that since Cocktail only supports pose maps, sketch maps, and segmentation maps, we only input pose maps and sketch maps when conducting experiments with Cocktail. Although the models of other approaches such as AnyControl [39] are public, their code cannot be successfully run after many attempts. More implementation details can be found in the supplementary material.
4.1 Main Results
Figure 6: Comparison of multi-condition image generation by PixelPonder with different combinations against single-condition image generation by ControlNet++ [20].
@@ 4.1 Main Results
MultiGen-20M Subject-200K
Methods Conditions FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
Source - - 80.69 - 69.30 - 76.48 - 66.96
Controlnet++ [20] canny 17.69 80.64 36.69 65.67 20.89 79.42 43.57 67.07
Controlnet++ [20] hed 13.93 81.07 42.12 71.22 15.67 78.37 56.21 68.27
Controlnet++ [20] depth 17.56 81.12 27.79 71.23 24.90 80.01 32.70 69.75
Ours canny+depth 15.99 76.52 38.53 68.26 17.04 77.46 55.67 67.22
Ours canny+hed 17.69 76.31 38.23 67.14 16.69 77.26 55.20 66.19
Ours canny+hed+depth 16.55 76.58 38.59 67.58 16.02 77.41 55.49 66.43
Ours all 11.85 78.60 43.99 69.54 10.61 77.12 66.60 67.32

### v2

# Yanjie Pan1∗, Qingdong He2∗, Zhengkai Jiang3, Pengcheng Xu4, Chaoyi Wang5, Jinlong Peng2, Haoxuan Wang1, Yun Cao2, Zhenye Gan2, Mingmin Chi1, Bo Peng1, Yabiao Wang2

@@ Abstract
Recent advances in diffusion-based text-to-image generation have demonstrated promising results through visual condition control. However, existing ControlNet-like methods struggle with compositional visual conditioning - simultaneously preserving semantic fidelity across multiple heterogeneous control signals while maintaining high visual quality, where they employ separate control branches that often introduce redundant guidance during the denoising process, leading to structural distortions and artifacts in generated images. To address this issue, we present PixelPonder, a novel unified control framework, which allows for effective control of multiple visual conditions under a single control structure. Specifically, we design a patch-level adaptive condition selection mechanism that dynamically prioritizes spatially relevant control signals at the sub-region level, enabling precise local guidance without global interference. Additionally, a time-aware control injection scheme is deployed to modulate condition influence according to denoising timesteps, progressively transitioning from structural preservation to texture refinement and utilizing the control information from different categories to promote finer image generation. Extensive experiments demonstrate that PixelPonder surpasses previous methods across different benchmark datasets, showing superior improvement in spatial alignment accuracy while maintaining high textual semantic consistency.
@@
Figure 1: Visualization comparison between Uni-ControlNet (Zhao et al. 2024) and our proposed method in different conditional controls with the same text prompt. (a, left two columns) Text and various visual controls, where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ denotes canny, sketch, depth and pose map respectively. (b, middle three or four columns) Generation results from UniControlNet. (c, last column) Generation results from our PixelPonder. Previous methods struggled to generate coherent results under multiple conditions, while our results maintain strong similarity to the respective visual controls.
Introduction
@@
The emergence of diffusion-based text-to-image synthesis (Rombach et al. 2022; Dhariwal and Nichol 2021; Peebles and Xie 2023; Podell et al. 2023; Ramesh et al. 2021, 2022) have driven significant progress in single-modality conditioned generation frameworks (Ye et al. 2023; Li et al. 2025; Mou et al. 2024; Cao et al. 2024; Li et al. 2023). While these approaches demonstrate enhanced controllability through visual-textual alignment, fundamental limitations persist due to modality-specific biases in visual conditioning. As evidenced in (Si et al. 2024; He et al. 2024), distinct visual control signals exhibit complementary characteristics, such as depth maps effectively regulate inter-object spatial relationships but lack fine-grained object details, whereas canny maps capture precise texture contours while disregarding global structural context. This modality specialization gap raises a critical challenge: How to establish an effective fusion paradigm for multi-visual condition control that synergistically integrates heterogeneous visual features while resolving inter-modality conflicts?
@@
In practical applications, it is often necessary to describe the visual features of a significant object using multiple visual conditions to achieve precise control over its generation. Users typically desire to simultaneously control both the layout and the intricate details. However, visual conditions that are rich in layout and detail information are challenging to integrate into a single visual condition map. A common scenario involves using canny maps to represent details and depth maps to control the layout. In such case, these visual conditions are highly overlapping and complementary to one another.
@@
Bridging the redundant information among these visual conditions and reasonably utilizing the emphasis of various visual conditions to coordinate the generation of the object is exactly the issue that this paper wants to focus on. As shown in Fig.2(a), prior efforts including Cocktail (Hu et al. 2023) and Uni-ControlNet (Zhao et al. 2024) propose to consolidate various visual conditions into unified latent representations. As illustrated in Fig.1, generating normal images under multi-modal and coupled visual condition controls remains a difficult task for these approaches. While this coarse-grained unification alleviates inter-condition conflicts through dimensionality reduction, it remains suboptimal in preserving modality-specific attributes critical for precise control (e.g., high-frequency textures in cannys vs. low-frequency structures in depth). The inherent information bottleneck induced by latent compression disproportionately attenuates complementary cross-modal features, fundamentally limiting their capacity to synergize discriminative visual cues. This constitutes a critical limitation: current unification paradigms prioritize conflict mitigation over inter-modal feature reinforcement.
@@
Figure 2: Comparison of different schemes dealing with multiple condition. (a) Existing methods integrate various visual controls into a highly consistent visual control signal in the temporal domain through concatenation or weighted summation. (b) We employ dynamic patch selection to reconstruct visual conditions at the patch level, enabling multi-visual controllable generation in the generative model.
@@
To address the aforementioned challenges, we introduce an innovative visual condition combination method termed PixelPonder. As illustrated in Fig.2(b), our method distinguishes itself from previous approaches by integrating diverse visual conditions at the patch level. Recognizing that visual conditions enriched with low-frequency or high-frequency signals significantly influence various denoising stages during the image generation process (Ho, Jain, and Abbeel 2020), we have developed an adaptive patch selection mechanism that dynamically adjusts with timesteps. This mechanism decouples each visual condition in the temporal domain, enabling the use of distinct combinations of visual conditions to regulate the denoising process at different stages. Our proposed combination method amalgamates various visual conditions with greater granularity, and facilitating complementary control of high- and low-frequency information within the temporal domain.
It is worth noting that when operating under a single visual condition control, our proposed method effectively reverts to the single visual control scheme. The efficacy of ControlNet’s single visual condition control has been extensively validated in numerous studies (Zhang, Rao, and Agrawala 2023; Zhao et al. 2024; Qin et al. 2023). Therefore, this paper does not investigate the generative capabilities of our method under single visual condition scenarios. Our main contributions are summarized as follows:
@@
  • Flexible Condition Combination: Our PixelPonder supports flexible visual control combination, addressing the redundant information conflict of multiple visual conditions for the same object.
@@
Related Work
Diffusion-based Generative Models
@@
Diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Nichol et al. 2021; Podell et al. 2023; Ramesh et al. 2021; Chen et al. 2023) have made remarkable strides in the field of image generation. Thanks to improvements in sampling strategies (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Song et al. 2020), as well as the application of latent diffusion methods (Rombach et al. 2022), the computational resources required for these models to generate high-quality images have been substantially reduced, demonstrating promising prospects in various applications (Cao et al. 2024; Brooks, Holynski, and Efros 2023; Hertz et al. 2022; Ramesh et al. 2022). In the domain of text-to-image generation, diffusion models (Rombach et al. 2022) combined with U-Net (Ronneberger, Fischer, and Brox 2015) have successfully integrated text control into the image generation process through cross-attention mechanisms. To further enhance generative capabilities, the Transformer architecture has been integrated into diffusion models (Peebles and Xie 2023; Chen et al. 2024; Esser et al. 2024). Building on this foundation, FLUX (Labs. 2024), which introduces flow matching objectives, has achieved state-of-the-art performance in image generation.
@@
Figure 3: Overall pipeline of the proposed. The Patch Adaptaiom Module segments various visual conditions $\{V_{1},V_{2},\ldots,V_{n}\}$ into patches and dynamiclly reorganizes the visual conditions at the patch level to obtain a unified visual control signal $SP_{out}$. Subsequently, $SP_{out}$ is sent to the control network, with the Time-aware Control Injection scheme to achieve high-quality image generation under multiple visual conditions.
Controllable Text-to-Image Diffusion Models
@@
The generation of images controlled by textual descriptions often lacks spatial control. To enhance spatial control capabilities, numerous studies (Mou et al. 2024; Zhang, Rao, and Agrawala 2023; Li et al. 2023; Huang et al. 2023; Cheng et al. 2023; Yang et al. 2023a; Zhang et al. 2023) have focused on incorporating visual control signals into the generation process. In this context, ControlNet (Zhang, Rao, and Agrawala 2023) proposed a spatial control network architecture, encoding additional visual conditions into latent representations and injecting control signals into the corresponding backbone network using zero convolution, thus achieving a spatial alignment mechanism within diffusion models. Due to its straightforward approach and effective control results, it has been widely adopted in various text-to-image diffusion models (Rombach et al. 2022; Labs. 2024). Furthermore, recent research has used prompt engineering techniques (Li et al. 2023; Yang et al. 2023a) and inter-attention constraints (Chen, Laina, and Vedaldi 2024; Xie et al. 2023) to facilitate more orderly generation.
@@
Some studies (Qin et al. 2023; Mou et al. 2024; Wang et al. 2024; Bar-Tal et al. 2023) have addressed the differences and diversity of various visual conditions in different modalities using the mix of experts (MOE) to decouple different visual conditions into distinct feature subspaces, thus establishing a unified visual control architecture. Subsequently, other research has explored the possibility of generating multiple condition controls within a single diffusion model through simple feature aggregation methods (Hu et al. 2023; Zhao et al. 2024). However, these methods are based on the low coupling of various conditions, which limits their effectiveness in tasks requiring the control of complementary conditions.
Method
Preliminary
@@
Diffusion Transformer (DiT)(Vaswani 2017; Peebles and Xie 2023) is applied in models such as FLUX  (Labs. 2024) and Stable Diffusion 3 (Esser et al. 2024). Compared to the Latent Diffusion Model (Rombach et al. 2022), these models utilize a transformer (Ruan et al. 2023) instead of U-Net as the denoising network to iteratively process noise tokens.
@@
A model with a DiT architecture handles noisy image tokens $X\in\mathbb{R}^{N\times d}$ and text condition tokens $Y\in\mathbb{R}^{M\times d}$, where $d$ is the dimension of the embedding layer, $N$ and $M$ represent the number of tokens for images and text, respectively. Additionally, they redefine the mapping from images to noise using flow matching. In this rectified mapping, the data at time $t$, denoted as $z_{t}$, is transformed from the real image $x_{0}\in\mathbb{R}^{C\times H\times W}$ to pure noise $N\sim\mathcal{N}(0,I)$ at the following position:
@@
ControlNet (Zhang, Rao, and Agrawala 2023) is designed to inject visual control information into a denoising network, thereby guiding the iterative process of noise tokens. Specifically, ControlNet freezes the parameters $\Theta$ of the backbone network $F(\cdot;\Theta)$ and creates a trainable copy $\Theta_{c}$ of these parameters to incorporate zero-initialized convolutional layers $Z$. When applied to generative models based on the U-Net architecture, a common practice is to add the visual condition $C_{i}$ to the noise $x_{t}$ as input to the trainable copy, while utilizing other signals $s$ that is consistent with the backbone network. This process can be formalized as follows:
@@
where $O_{t}$ refers to the time step information, and its specific role will be discussed in next subsection.
Assuming that the visual input to the PAM consists of $N$ types of condition images, denoted as $V_{1},V_{2},\ldots,V_{n}$, we use the corresponding visual condition feature encoders $E_{i}$ to obtain various visual condition features $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$, $C_{k}\in\mathbb{R}^{e\times d}$ and $e$ denotes the feature dimension. To analyze and process the visual condition features in a more fine-grained manner, we can consider the visual condition feature $C_{k}$ as represented by patches $P_{i}^{m}\in\mathbb{R}^{d}$, expressed as:
$C_{k}=[P_{i}^{0},\ldots,P_{i}^{m}],$ (5)
@@
[…]
@@
7:  $M_{\text{prob}}=[\text{ISB}_{\text{C}}(\mathcal{C}_{i},O_{t},Y)]+\text{ISB}_{\text{SP}}(SP_{i})$
8:  $M_{N,i}=\text{Zero}(\text{Top}_{r}(M_{prob}))$, $M_{sp,i}=\sim M_{N,i}$
9:  $\mathcal{C}_{i+1}=\mathcal{C}_{i}M_{N,i}$
10:  $SP_{i+1}=SP_{i}+M_{sp,i}\mathcal{C}_{i}$
11:  $i\leftarrow i+1$
12:end while
13:return $SP_{i}$
@@
Specifically, to achieve $F_{\text{sp}}$, we design the Image Stream Block (ISB). To ensure that the textual control signal remains consistent globally, We modified the DSB text input selection strategy to adopt a globally unified textual feature representation $Y$ as the input textual information. The ISB is used to obtain the weights of each patch in various visual condition features under the current textual control $Y$ at time step $t$, denoted as $\mathcal{W}_{c}=\{W_{0,i},\ldots,W_{n,i}\}$. Furthermore, to meet requirement b) of $F_{\text{sp}}$, we apply an ISB to $SP_{i}$ outside of $\mathcal{C}_{i}$ to obtain a global weight $W_{\text{sp}}$ that represents the global composition information of $SP_{i}$ under the current $U_{i}$. The final obtained patch selection probability map $M_{\text{prob}}$ is given by:
@@
Figure 4: Denoising results at steps 25, 20, 15, 10, 5, and 1, where each pair includes the generated result and the selection results $SP_{out}$ at the current time step. In each $SP_{out}$, canny, pose, depth and sketch are indicated in red, blue, green, and yellow respectively.
Fig.4 illustrates the visualization results of $SP_{out}$ obtained by PAM in different time steps, as well as the generated images. In the denoising process, the generated images exhibit a noticeable tendency towards patchiness. The segmentation contours between the generated image patches are prominent. Both examples demonstrate the characteristics of PAM: a) PAM exhibits a balanced tendency in its selections over various time steps. b) For patches at the same location, the global selections include patches from all categories.
@@
Time-aware Control Injection
@@
Methods Conditions FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) Inference Time ($\downarrow$)
T2I-Adapter all 66.95 71.47 24.03 57.95 64.72 74.35 31.76 55.07 9.33
Uni-ControlNet all 32.58 78.08 29.37 65.85 44.35 77.40 37.98 66.75 7.22
UniControl all 25.15 74.09 35.58 72.05 30.95 72.96 47.31 67.50 7.39
Cocktail pose+hed 24.67 79.87 24.14 67.13 51.51 79.38 29.59 64.16 4.94
PixelPonder(Ours) all 11.85 78.60 43.99 69.54 10.61 77.12 66.60 67.32 9.59
Table 1: Comparison on the MultiGen-20M (Li et al. 2025) and Subject-200K (Tan et al. 2024) datasets. The best and second-best results are highlighted in bold and underline, respectively. In the Conditions column, “all” denotes the use of canny maps, sketches, pose maps, and depth maps simultaneously during generation.
We validate the effectiveness of PixelPonder on the MultiGen-20M (Li et al. 2025) and Subject-200K(Tan et al. 2024) datasets. Our evaluation primarily focuses on several leading methods, including T2I-Adapter (Mou et al. 2024), Uni-ControlNet (Zhao et al. 2024), UniControl (Qin et al. 2023), and Cocktail (Hu et al. 2023) in the realm of multiple controllable text-to-image diffusion models. It is worth noting that we conducted experiments on Cocktail using only pose and sketch maps, as it only supports pose, sketch, and segmentation maps. Although the models of other approaches such as AnyControl (Sun et al. 2024) are public, their code cannot be successfully run after many attempts. More implementation details can be found in the supplementary material.
Main Results
@@
Comparison of CLIP Score. Tab.1 compares CLIP Score metric across methods under maximal conditions. The results reveal a consistent trade-off: methods with fewer visual constraints achieve higher text adherence. Specifically, PixelPonder demonstrates superior performance under full conditional control, outperforming the compared methods. Although Cocktail attains higher scores under significantly fewer visual constraints, the strong text adherence of Cocktail significantly compromises the consistency of its visual performance. As illustrated in Fig.5, the images generated by Cocktail exhibit pronounced issues of object loss and mutations when compared to those produced by other methods. This highlights PixelPonder’s unique capability to balance precise visual control with competitive text adherence.
Figure 6: Comparison of multi-condition image generation by PixelPonder with different combinations against single-condition image generation by ControlNet++.
Comparison of Inference Speed. Tab.1 presents the inference speed. Due to the fact that Cocktail involves only half the number of conditions compared to other methods, any direct comparison of time efficiency may not be entirely representative or meaningful. Patch Adaptation introduces a marginal increase in generation time; however, given the substantial improvement in image quality, the associated computational overhead is acceptable.
Condition Types Comparison. To examine the effect of the number of condition types, Fig.6 shows the visual results under various condition combinations, with qualitative evaluation metrics provided in Tab.2. Examples in Fig.6 demonstrate that as the number of conditions increases, the layout and texture within images become increasingly refined and accurate, demonstrating that the control effects of different visual conditions are not identical. Furthermore, as can be seen from Tab.2, an increase in visual conditions is beneficial for the overall quality and controllability of the images, which aligns with the observed visual effects. As evidenced by the comparison with ControlNet++ (Li et al. 2025), integrating multi-conditions improves control precision and enables finer-grained manipulation of the generated output.
MultiGen-20M
Methods Conditions FID ($\downarrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
Source - - - 69.30
ControlNet++ canny 17.69 36.69 65.67
ControlNet++ hed 13.93 42.12 71.22
ControlNet++ depth 17.56 27.79 71.23
Ours pose 30.00 27.91 58.16
Ours pose+depth 19.76 31.63 62.63
Ours pose+depth+hed 12.41 40.48 67.82
Ours all 11.85 43.99 69.54